# Filtered Direct Preference Optimization

### v2

@@ 5.1 Pretrained Models
In our experiments, we employed pretrained LMs provided in the Pythia suite by Biderman et al. (2023) of two different sizes: 1.4B and 160M models. In Section 5.4, we present the primary findings of our study using a 1.4B-sized LM to highlight the effectiveness of fDPO in an LLM. Due to computational resource constraints, a comprehensive examination of the 160M LM is provided in Section 5.5. In the preliminary setup, each LM was subjected to SFT using the demonstration dataset from AlpacaFarm (Dubois et al., 2023). These prepared LMs, denoted as $\pi_{\theta}$, were then used as the initial LMs for our experiments.
For the (proxy) RM, we used Pythia models of varying sizes: 14M, 70M, and 160M models. To circumvent the high costs associated with human evaluation, similar to other studies Dubois et al. (2023); Rafailov et al. (2023), we utilized a large-scale human preference model as the gold RM. Specifically, ”OpenAssistant/reward-model-deberta-v3-large-v2”  [^3] https://huggingface.co/OpenAssistant/reward-model-deberta-v3-large-v2 model was employed for this purpose.
5.2 Data
We utilized the AlpacaFarm instruction dataset (Dubois et al., 2023) for our experiments. The dataset consisted of 20,000 training samples. For evaluation, a random selection of 200 samples from in a pool of 2000 was used in each independent run to ensure consistency and unbiased results.
@@ 5.2 Data
Given our focus on dataset quality, we employed gold RM and BoN sampling (Stiennon et al., 2020; Nakano et al., 2021) to create three types of pairwise preference datasets:
@@
This dataset was created using the conventional manner. For each prompt $x$, the LM $\pi_{\theta}$ generated two responses. These responses were then evaluated by the gold RM, with the higher-scoring response designated as $y_{c}$ and the other as $y_{r}$. This formed the preference dataset $\mathcal{D}$ samples $(x,y_{c},y_{r})$. For brevity, this dataset is referred to as the ‘low dataset’.
@@
Adopting the approach from (Pace et al., 2024), we used BoN sampling to create responses of higher quality. Specifically, for each prompt $x$, the LM $\pi_{\theta}$ generated $16$ responses. These responses were then evaluated by the gold RM, and the highest-scoring response was selected as $y_{c}$, with one randomly selected from the remaining 15 responses labeled as $y_{r}$. Due to the probabilistic nature of outputs of $\pi_{\theta}$, this approach is likely to yield $y_{c}$ responses of higher quality (as indicated by gold RM scores) compared to the $y_{c}$ responses in the low-quality dataset. For simplicity, this dataset is referred to as the ‘high dataset’.
@@
This dataset was created by mixing the low-quality and high-quality datasets in a 50/50 ratio, ensuring no overlap in prompts between the two. This dataset is referred to as the ‘mix dataset’.
We provide the evaluation scores of the gold RM for these datasets in Table 3 in Appendix B.
@@ 5.3 Effect of Data Quality to Performance of RLHF and DPO
Our preliminary experiment investigates the sensitivities of (RM-based) RLHF and (RM-free) DPO to the quality of the datasets employed with the 160M-sized LM. Here, we used the high-quality and mixed-quality datasets. We used an RLHF implementation in the TRL library, in which a 70M-sized RM is trained and then used for RL fine-tuning. The evaluation is based on three independent runs, with each run employing different evaluation prompts that could vary ranges of the average of gold rewards. For a coherent comparison, gold rewards are normalized, setting the initial performance of LM to $0.0$ and post-training performance with RLHF or DPO to $1.0$.
Figure 1 (A) shows the results, where the mean and standard error of the normalized gold reward with three independent runs are presented. Notably, while DPO experienced a decline in efficacy when trained on the mixed-quality dataset relative to the high-quality one, RLHF showed an intriguing resilience, sustaining comparable performance levels across both datasets. This differential impact starkly highlights the greater susceptibility of DPO to dataset quality, suggesting that the RM-based approach, including fDPO, may offer more stable performance when the preference dataset quality cannot be consistently assured. All experiments are conducted using a single NVIDIA A100 accelerator. Further details of the experimental parameters are provided in Table 1 in Appendix B.
5.4 Evaluation of fDPO
We evaluate fDPO and DPO when trained using a 1.4B-sized LM $\pi_{\theta}$ on the mixed-quality dataset, where fDPO used a 160M-sized RM that was trained with the same dataset. The evaluation is based on three independent runs, where the gold reward was normalized with the performance of DPO with high-quality dataset, as in Section 5.3. The epoch number for DPO was set to 5, which avoided overoptimization while ensuring the learning convergence. In the case of fDPO, we adapted the epoch count to double that of DPO, up to 10 epochs, or ceased learning when the remaining sample size was reduced to less than 1% of the original due to the filtering process, meaning the maximum filtered rate $\gamma=0.99$.
Figure 2 present the results of DPO and fDPO. The results shows that the performance of DPO trained on the high-quality dataset and fDPO trained on the mixed-quality dataset were on par. It indicates that fDPO has successfully circumvented the performance decline typically observed with DPO, thereby showcasing its potential to improve DPO performance where dataset quality is inconsistent. Corresponding learning curves are included in Appendix B.
@@ 5.4 Evaluation of fDPO
5.5 Detailed Experiments
We delve into an extensive analysis of fDPO using a 160M LM. We set the number of epochs to $8$ for DPO to ensure convergence, resulting in a maximum of $16$ epochs for fDPO. We maintained the maximum filtered rate ($\gamma=0.99$) as unchanged.
Performance comparison with DPO.
Figure 1 (B) illustrates the performances of LMs trained with DPO and fDPO using the mixed-quality dataset. The results are consistent with those obtained from the larger 1.4B-sized LM, reaffirming the advantage of fDPO with the mixed-quality dataset. Additionally, we conducted an experiment using only a low-quality dataset, which revealed a surprising improvement of 7.42% (standard error: 1.89%) despite the presumed uniformity of response quality. This improvement suggests it effectively discriminates subtle quality variations, enhancing overall performance by eliminating less optimal data, even within uniformly labeled datasets.
Analysis of configuration parameters.
@@ 5.5 Detailed Experiments
Reducing randomness of LMs was hypothesized to enhance fDPO’s performance by minimizing the variance in quality of the LM-generated responses. However, as Figure 5 indicates, reducing randomness did not yield improvements, and in some cases, it led to worse performance. This outcome may be attributed to a discrepancy between inference-time and training-time randomness.
@@
Finally, we explored different criteria for discarding data. As stated in line 8 of Algorithm 1, the original criterion was discarding a sample even if the reward of the LM-generated response $y$ is only marginally higher than that of $y_{c}$ in the dataset. Considering potential errors in proxy rewards and the probabilistic nature of LMs, we introduced a margin $\epsilon$ to the discarding criterion: $r(x,y)>r(x,y_{c})+\epsilon$. Figure 5 presents the results, showing that a certain margin size can indeed be effective in enhancing fDPO’s performance.
Analysis of filtered samples.
We examined how data was selectively discarded throughout the learning process of fDPO with the mix-quality dataset. Figure 6 presents the unfiltered ratio, accuracy, precision, and recall at each epoch. The unfiltered ratio reflects the proportion of data that remains after filtering. Accuracy reflects the overall correctness of the filtering decisions, both for deletion and retention of samples, based on their gold reward quality. Precision measures how accurately the samples decided for deletion were actually of lower quality, while recall evaluates the success in identifying and discarding all samples that warranted removal. The result of the unfiltered ratio indicates an exponential decay in the number of samples used in each epoch. The consistency of accuracy and precision across epochs suggests that data was discarded with a constant efficiency. The lower precision compared to accuracy can be attributed to the relatively small number of samples that warranted removal. Conversely, recall decreases with progressing epochs. This decline can be tied to the static errors within the proxy RM, leading to consistently overestimated $y_{c}$ samples, thus increasing their relative proportion over time. The figure contrasts various margin settings with the no-margin condition ($\epsilon=0$), revealing that larger margins lead to slower filtering speeds. Notably, as the margin increases, precision improves at the expense of recall. This trade-off underscores the importance of carefully tuning the margin parameter $\epsilon$ to balance filtering efficacy.
Figure 6: Unfiltered ratio, accuracy, precision, and recall throughout epochs in fDPO, comparing the effects of no margin ($\epsilon=0$) with various margin levels ($\epsilon>0$) on the filtering condition.
@@ 6 Conclusions
This study explores how the quality of a preference dataset impacts language models (LMs) optimized using direct preference optimization (DPO), especially when compared with an ordinary reinforcement learning from human feedback (RLHF) method. We discovered that the impact of dataset quality is significantly more pronounced in DPO. To address the challenges posed by low-quality data within DPO, we proposed filtered DPO (fDPO). This approach utilizes a reward model (RM) to identify and discard lower-quality data, effectively refining the DPO process. Our experiments demonstrated that fDPO improved DPO’s performance across different LM sizes, effectively handling datasets with quality discrepancies. This finding highlights the practicality of fDPO, particularly for applications where data quality can vary or is challenging to control. Intriguingly, fDPO also presented potential performance gains in datasets with standard quality variance.
Several avenues for future research emerge from this work. One vital area involves expanding our experimentation to diverse, real-world preference datasets. In this study, we employed artificially created datasets with varied response qualities. However, merging multiple existing preference datasets, which naturally contain varied qualities, could offer more realistic insights into fDPO’s applicability and robustness. Exploring how fDPO performs with these organically varied datasets will be crucial for understanding its effectiveness in more complex, real-application scenarios.
Further, the uncertainties associated with proxy RM estimation and response generation of LM require more in-depth exploration. Our initial assessments suggest robustness against these factors; however, a comprehensive investigation is essential to validate these findings. Additionally, while fDPO currently focuses solely on the quality of chosen responses, considering rejected responses presents an exciting direction for future work.
Appendix A Appendix
A.1 Sensitivity analysis of chosen and rejected responses
This subsection presents a theoretical analysis focused on the differential sensitivity of the DPO algorithm to chosen ($y_{c}$) and rejected ($y_{r}$) responses. The analysis elucidates how the DPO’s update gradient affects the probability of chosen responses relative to rejected ones, which is a key consideration in designing our proposed approach, fDPO. This understanding is vital to enhance the efficiency of DPO, which fDPO achieves by selectively discarding low-quality $y_{c}$ samples during training.
For simplicity in the following analysis, we will occasionally omit the prompt $x$ and denote $\pi_{\theta}(y\,|\,x)$ simply as $\pi_{\theta}(y)$.
Proposition 1.
@@ A.1 Sensitivity analysis of chosen and rejected responses
$|||\nabla_{\theta}\log\pi_{\theta}(y_{c})||-||\nabla_{\theta}\log\pi_{\theta}(y_{r})|||\leq o(||\nabla_{\theta}\log\pi_{\theta}(y)||),\quad\forall y\in\{y_{c},y_{r}\}$
$|\nabla_{\theta}\log\pi_{\theta}(y_{c})^{\!\top}\nabla_{\theta}\log\pi_{\theta}(y_{r})|^{\frac{1}{2}}\leq o(||\nabla_{\theta}\log\pi_{\theta}(y)||),\quad\forall y\in\{y_{c},y_{r}\}$
$\pi_{\theta}(y_{c})/\pi_{\theta}(y_{r})=\delta.$
When, following the DPO algorithm, the parameter $\theta$ is updated using the gradient given by
$w(y_{c},y_{r})(\nabla_{\theta}\log\pi_{\theta}(y_{c})-\nabla_{\theta}\log\pi_{\theta}(y_{r})),$
the sensitivity of $\pi_{\theta}(y_{c})$, defined as the magnitude of change in probability, $\Delta\pi_{\theta}(y)$, is approximately $\delta$ times higher than that of $\pi_{\theta}(y_{r})$.
@@
We notate the DPO gradient as $\Delta\theta=w(y_{c},y_{r})(\nabla_{\theta}\log\pi_{\theta}(y_{c})-\nabla_{\theta}\log\pi_{\theta}(y_{r}))$. The DPO gradient $\Delta\theta$ directly influences the rate of change in the probabilities $\pi_{\theta}(y_{c})$ and $\pi_{\theta}(y_{r})$. The change in probabilities can be approximated as
@@
$\simeq\pi_{\theta}(y)\Delta\theta^{\!\top}\nabla_{\theta}\log\pi_{\theta}(y)$
Given the assumptions, the magnitudes of the gradients for $\log\pi_{\theta}(y_{c})$ and $\log\pi_{\theta}(y_{r})$ are similar and their inner product is small, indicating that the direction of their gradients is nearly orthogonal. Hence, the impact of $\Delta\theta$ on $\log\pi_{\theta}(y_{c})$ and $\log\pi_{\theta}(y_{r})$ would be similar in magnitude but differ in direction. However, due to the ratio $\pi_{\theta}(y_{c})/\pi_{\theta}(y_{r})=\delta$, the rate of change in $\pi_{\theta}(y_{c})$ is amplified by a factor of $\delta$ compared to $\pi_{\theta}(y_{r})$. Thus, under the DPO update, $\pi_{\theta}(y_{c})$ demonstrates a sensitivity approximately $\delta$ times higher than that of $\pi_{\theta}(y_{r})$. ∎
As the training progresses in DPO, it is generally observed that the ratio $\delta$, representing the likelihood of $y_{c}$ over $y_{r}$, tends to exceed 1. This phenomenon indicates an increased sensitivity towards the chosen responses, emphasizing the criticality of their quality within the DPO framework. Consequently, the presence of low-quality chosen responses in the dataset can significantly impede the effectiveness of DPO.
@@
Appendix B Details of experiments
B.1 hyperparameters
We provide details of the hyperparameters used in our experiments. Table 1 presents the training parameters for DPO and RLHF, while Table 2 illustrates the parameters for response generation.
Table 1: Experimental Parameters of DPO and PPO
Parameter DPO (160M) DPO (1.4B) RLHF (160M)
epochs 8 (16 for fDPO) 5 (10 for fDPO) 3
batch size 32 16 16
gradient accumulation steps 2 4 2
learning_rate $1\times 10^{-6}$ $2\times 10^{-7}$ $1\times 10^{-6}$
$\beta$ 0.1 0.1 0.1
PPO epochs - - 4
number of rollouts - - 128
Table 2: Response Generation Params
Parameter value
Top p 0.9 (1.0 for fDPO and PPO)
Top k 0
Temperature 1.0
Max response tokens 64
B.2 Additional results
The evaluation scores of the gold reward models for each preference dataset (’high’, ’low’, ’mix’) are detailed in Table 3.
Table 3: The evaluation scores of gold rewards for each dataset
Model Size Quality Chosen Mean Rejected Mean Overall Mean

### v3

@@ 5.1 Pretrained Models
We employed pretrained LMs provided in the Pythia suite by Biderman et al. (2023) of two different sizes: 1.4B and 160M models, in experiments on the AlpacaFarm dataset. In experiments on the Anthropic HH datasets, we used the 2.8B-sized Pythia model. Due to computational resource constraints, a comprehensive examination of the 160M LM is provided in Sections 5.3 and 5.4. In the preliminary setup, each LM was subjected to SFT using the demonstration data in the AlpacaFarm dataset or the chosen responses from the preference data in the Anthropic HH datasets, as the Anthropic HH datasets do not contain demonstration data. These prepared SFT models, denoted as $\pi_{\theta}$, were then used as the initial LMs for our experiments.
For the (proxy) RM, we used Pythia models of varying sizes: 14M, 70M, and 160M models, with 160M being the default unless otherwise specified. To circumvent the high costs associated with human evaluation, similar to other studies Dubois et al. (2023); Rafailov et al. (2023), we utilized a large-scale human preference model as the gold RM. Specifically, “OpenAssistant/reward-model-deberta-v3-large-v2” [^3] https://huggingface.co/OpenAssistant/reward-model-deberta-v3-large-v2 model was employed for this purpose. We adjusted the reward zero point such that the average reward of the initial LM (SFT model) is set to zero. Additionally, in Section 5.5, we employed GPT-4o for evaluation as an alternative to human assessment, accessed via Azure OpenAI. [^4] https://learn.microsoft.com/en-us/azure/ai-services/openai/concepts/models The specific model used was “gpt-4o” with the version “2024-05-13”.
5.2 Datasets
We used the AlpacaFarm dataset (Dubois et al., 2023) and the Anthropic HH datasets Bai et al. (2022). The AlpacaFarm dataset consists of 169,352 demonstration (SFT) samples, 20,000 training samples, and 2,000 test samples. The Anthropic HH datasets include two subtypes of datasets: helpfulness and harmlessness datasets. The former consists of 43,835 training samples and 2,354 test samples. The latter consists of 42,537 training samples and 2,312 test samples.
@@ 5.2 Datasets
Given our focus on dataset quality, in experiments on the AlpacaFarm dataset, we employed gold RM and BoN sampling (Stiennon et al., 2020; Nakano et al., 2021) to create three types of pairwise preference datasets:
@@
This dataset was created using the conventional manner. For each prompt $x$, the LM $\pi_{\theta}$ generated two responses. These responses were then evaluated by the gold RM, with the higher-scoring response designated as $y_{c}$ and the other as $y_{r}$. This formed the preference dataset $\mathcal{D}$ samples $(x,y_{c},y_{r})$. For brevity, this dataset is referred to as the low dataset.
@@
Adopting the approach from (Pace et al., 2024), we used BoN sampling to create responses of higher quality. Specifically, for each prompt $x$, the LM $\pi_{\theta}$ generated $16$ responses. These responses were then evaluated by the gold RM, and the highest-scoring response was selected as $y_{c}$, with one randomly selected from the remaining 15 responses labeled as $y_{r}$. Due to the probabilistic nature of outputs of $\pi_{\theta}$, this approach is likely to yield $y_{c}$ responses of higher quality (as indicated by gold RM scores) compared to the $y_{c}$ responses in the low-quality dataset. For simplicity, this dataset is referred to as the high dataset.
@@
This dataset was created by mixing the low-quality and high-quality datasets in a 50/50 ratio, ensuring no overlap in prompts between the two. This dataset is referred to as the mix dataset.
We provide the evaluation scores of the gold RM for these datasets in Table 4 in Appendix C.
For experiments on the Anthropic HH datasets, we created mix-quality datasets by combining original responses from the dataset and those generated by the SFT model. Details are provided in Section 5.5.
@@ 5.3 Effect of Data Quality to Performance of RLHF and DPO
Our preliminary experiment investigates the sensitivities of (RM-based) RLHF and (RM-free) DPO to the quality of the datasets employed with the 160M-sized LM. Here, we used the high-quality and mixed-quality datasets. For RLHF, the 70M-sized RM was trained from the same datasets and used for RL fine-tuning with PPO. The evaluation is based on five independent runs.
Figure 1 (A) shows the results, where the mean and standard error of the gold reward with five independent runs are presented. Notably, while DPO experienced a decline in efficacy when trained on the mixed-quality dataset relative to the high-quality one, RLHF showed an intriguing resilience, sustaining comparable performance levels across both datasets. This differential impact starkly highlights the greater susceptibility of DPO to dataset quality, suggesting that the RM-based approach, including fDPO, may offer more stable performance when the preference dataset quality cannot be consistently assured. However, RLHF’s overall gold reward was lower than DPO’s. Therefore, subsequent experiments focus on DPO.
5.4 Evaluation of fDPO on AlpacaFarm dataset
We evaluate fDPO and DPO when trained using a 1.4B-sized LM $\pi_{\theta}$ on the mixed-quality dataset, where fDPO used a 160M-sized RM that was trained with the same dataset. The evaluation is based on five independent runs. The epoch number for DPO was set to 5, which avoided overoptimization while ensuring the learning convergence. In the case of fDPO, we adapted the epoch count to double that of DPO, up to 10 epochs.
Figure 2 present the results of DPO and fDPO. The results shows that the performance of DPO trained on the high-quality dataset and fDPO trained on the mixed-quality dataset were on par. It indicates that fDPO has successfully circumvented the performance decline typically observed with DPO, thereby showcasing its potential to improve DPO performance where dataset quality is inconsistent. Corresponding learning curves are included in Appendix C.
@@ 5.4 Evaluation of fDPO on AlpacaFarm dataset
5.4.1 Detailed Evaluation
@@
We examines an extensive analysis of fDPO using a 160M LM. We set the number of epochs to $8$ for DPO to ensure convergence, resulting in a maximum of $16$ epochs for fDPO.
Performance comparison with DPO.
Figure 1 (B) illustrates the performances of LMs trained with DPO and fDPO using the mixed-quality dataset. The results are consistent with those obtained from the larger 1.4B-sized LM, reaffirming the advantage of fDPO with the mixed-quality dataset. Additionally, we conducted an experiment using only a low-quality dataset, which revealed a significant improvement of 4.10% (standard error: 1.87%) despite the presumed uniformity of response quality. This improvement suggests it effectively discriminates subtle quality variations, enhancing overall performance by eliminating less optimal data, even within uniformly labeled datasets.
Analysis of configuration parameters.
Dataset Method Gold RM Score (SFT=$0.0$) $\uparrow$ GPT-4o Evaluation (win rate vs. SFT) $\uparrow$
Helpful DPO 1.42 $\pm$ 0.08 0.543 $\pm$ 0.015
fDPO 1.94 $\pm$ 0.02 0.628 $\pm$ 0.001
Harmless DPO 2.66 $\pm$ 0.12 0.891 $\pm$ 0.003
fDPO 3.20 $\pm$ 0.06 0.944 $\pm$ 0.005
Table 1: Evaluation on the Anthropic HH datasets. The values represent the mean and standard error over 3 seeds.
@@
Reducing randomness of LMs during the filtering process was hypothesized to enhance fDPO’s performance by minimizing the variance in quality of the LM-generated responses used for filtering training samples. The idea was that more consistent response quality would lead to more reliable filtering decisions. However, as Figure 5 indicates, reducing randomness did not yield improvements, and in some cases, it led to worse performance. This outcome may be attributed to a discrepancy between inference-time and training-time randomness.
@@
Finally, we explored different criteria for discarding data. As stated in line 8 of Algorithm 1, the original criterion was discarding a sample even if the reward of the LM-generated response $y$ is only marginally higher than that of $y_{c}$ in the dataset. Considering potential errors in proxy rewards and the probabilistic nature of LMs, we introduced a margin $\epsilon$ to the discarding criterion: $r(x,y)>r(x,y_{c})+\epsilon$. Figure 5 presents the results, showing that larger margins generally lead to a decrease in performance, with the best results achieved when no margin is applied. This suggests that setting a margin $\epsilon$ is not necessary for enhancing fDPO’s performance. We further examined how samples were selectively discarded throughout the learning process of fDPO in Appendix C.3.2.
5.5 Evaluation of fDPO under Realistic RLHF Settings on Anthropic HH Datasets
We also conducted experiments on the Anthropic HH datasets, which consist of single-turn dialogues covering various topics such as academic questions or life guidance Bai et al. (2022). Here, we aimed to replicate a realistic RLHF setting where the number of high-quality responses created by humans is limited. Instead of generating all responses manually, SFT models are used to create response pairs, and human annotators only provide labels (chosen or rejected) to the pairs. This setup is cost-effective because generating high-quality responses manually is expensive, while annotating SFT-generated pairs is less so. This approach is consistent with the RLHF pipeline used in Ouyang et al. (2022); Pace et al. (2024); Yuan et al. (2024), which utilize unlabeled prompts effectively.
Specifically, we treated the original responses in the Anthropic HH datasets as high-quality responses, comprising 25% of the dataset. The remaining 75% of the responses were generated by the SFT model. These responses were then annotated as chosen or rejected by the gold RM.
The evaluation metrics used in this study included the gold RM score, as described in the previous sections, and an additional evaluation using GPT-4o to determine the win rate. The win rate indicates how often responses generated by the trained LM were preferred over those generated by the initial SFT model.
As shown in Table 1, based on three independent runs, fDPO outperformed the baseline in both evaluation metrics of the gold RM scores and GPT-4o win rates. The superior GPT-4o evaluation results suggest that fDPO is not merely optimizing for the reward model but is also learning to generate higher-quality responses from a human evaluation perspective. This demonstrates the effectiveness of our approach under realistic RLHF settings, providing a viable solution for scenarios where high-quality responses are limited.
@@ 6 Conclusions
This study explores how the quality of a preference dataset impacts LMs optimized using DPO, especially when compared with the RLHF method. We found that the quality of chosen responses significantly influences DPO performance. To address this, we proposed filtered DPO (fDPO), which uses a reward model to identify and discard lower-quality data, refining the DPO process. Our experiments demonstrated that fDPO improved DPO’s performance, effectively handling datasets with quality discrepancies. While the use of a reward model introduces additional computational costs and complexity, it allows for more effective leveraging of limited data. Overall, this highlights the practical value of fDPO’s approach, especially in scenarios where data quality is heterogeneous.
7 Limitations
The fDPO method shows promise, but it has some limitations. First, the method requires a reward model, which might be a drawback as it increases the complexity and computational time of the method. However, the availability of high-quality reward models provides an opportunity to leverage these high-end models within the DPO framework. Exploring the use of implicit rewards in DPO instead of an explicit reward model could also address some complications associated with training a separate reward model. Second, the algorithm is implemented in its simplest form, suggesting significant room for improvement and optimization. Third, our approach does not account for rejected responses, which could further enhance performance if considered. Finally, our experiments are limited to relatively small LLMs and comparisons with DPO. Future work should explore combining fDPO with other DPO-related extensions and conducting comparisons with other RLHF methods, especially with larger LLMs.
Appendix A Ethical considerations
This study addresses the challenge of aligning large language models with human preferences. We used publicly available datasets (AlpacaFarm and Anthropic HH), ensuring data transparency and privacy. While this study did not specifically evaluate models for biases, we acknowledge the significance of these considerations and commit to addressing them in future work.
Appendix B Justification on filtering chosen responses
To understand the impact of the quality of chosen responses on the performance of the DPO algorithm, we presents a theoretical analysis focused on the differential sensitivity of the DPO algorithm to chosen ($y_{c}$) and rejected ($y_{r}$) responses. The analysis elucidates how the DPO update affects the probability of chosen responses relative to rejected ones, which is a key consideration in designing our proposed approach fDPO. This understanding is vital to enhance the efficiency of DPO, which fDPO achieves by selectively discarding low-quality $y_{c}$ samples during training. For simplicity in this analysis, we will occasionally omit the prompt $x$, denoting $\pi_{\theta}(y\,|\,x)$ simply as $\pi_{\theta}(y)$.
Proposition B.1.
@@ Appendix B Justification on filtering chosen responses
  • the magnitudes of the gradients for $\log\pi_{\theta}(y_{c})$ and $\log\pi_{\theta}(y_{r})$ are similar, i.e.,
    $\|\nabla_{\theta}\log\pi_{\theta}(y_{c})\|\simeq\|\nabla_{\theta}\log\pi_{\theta}(y_{r})\|,$
  • the gradients for $\log\pi_{\theta}(y_{c})$ and $\log\pi_{\theta}(y_{r})$ are nearly orthogonal, i.e.,
    $\nabla_{\theta}\log\pi_{\theta}(y_{c})^{\!\top}\nabla_{\theta}\log\pi_{\theta}(y_{r})\simeq 0,$
  • the ratio of the probabilities is given by $\pi_{\theta}(y_{c})/\pi_{\theta}(y_{r})=\delta$.
When the DPO algorithm updates the parameter $\theta$ with
$\Delta\theta=\alpha\beta w(y_{c},y_{r})(\nabla_{\theta}\log\pi_{\theta}(y_{c})-\nabla_{\theta}\log\pi_{\theta}(y_{r})),$
where $\alpha$ is the learning rate and is sufficiently small, the sensitivity of $\pi_{\theta}(y_{c})$, defined as the magnitude of change in probability, $\Delta\pi_{\theta}(y)$, is approximately $\delta$ times higher than that of $\pi_{\theta}(y_{r})$.
@@
Since $\alpha$ is sufficiently small, which implies that the higher-order terms can be ignored, the variation in probabilities can be approximated as
@@
$\simeq\pi_{\theta}(y)\Delta\theta^{\!\top}\nabla_{\theta}\log\pi_{\theta}(y).$
Given the assumptions, the magnitudes of the gradients for $\log\pi_{\theta}(y_{c})$ and these gradients are nearly orthogonal. Hence, the impact of $\Delta\theta$ on $\log\pi_{\theta}(y_{c})$ and $\log\pi_{\theta}(y_{r})$ would be similar in magnitude but differ in direction. However, due to the ratio $\pi_{\theta}(y_{c})/\pi_{\theta}(y_{r})=\delta$, the rate of change in $\pi_{\theta}(y_{c})$ is amplified by a factor of $\delta$ compared to $\pi_{\theta}(y_{r})$. Thus, under the DPO update, $\pi_{\theta}(y_{c})$ demonstrates a sensitivity that is approximately $\delta$ times higher than that of $\pi_{\theta}(y_{r})$. ∎
As the training progresses in DPO, it is generally observed that the ratio $\delta=\pi_{\theta}(y_{c})/\pi_{\theta}(y_{r})$, representing how much more likely $y_{c}$ is compared to $y_{r}$, tends to exceed $1$. This phenomenon indicates an increased sensitivity towards the chosen responses, emphasizing the criticality of their quality within the DPO framework. Consequently, the presence of low-quality chosen responses in the dataset can significantly impede the effectiveness of DPO. Our proposed fDPO addresses this issue by selectively discarding samples with low-quality chosen responses during training, thereby enhancing the overall performance and robustness of the model.
@@
Appendix C Details of experiments
C.1 Hyperparameters
We provide details of the hyperparameters used in our experiments. The hyperparameters were chosen to optimize the performance of DPO. Table 2 presents the training parameters for DPO and RLHF of the AlpacaFarm dataset and DPO of the Anthropic HH dataset. Table 3 illustrates the parameters for response generation.
Parameter DPO 160M (AlpacaFarm) DPO 1.4B (AlpacaFarm) RLHF (AlpacaFarm) DPO 2.8B (Anthropic HH)
epochs 8 (16 for fDPO) 4 (8 for fDPO) 3 4 (8 for fDPO)
batch size 64 64 32 64
learning_rate $1\times 10^{-6}$ $2\times 10^{-7}$ $1\times 10^{-6}$ $5\times 10^{-6}$
$\beta$ 0.1 0.1 0.1 0.1
PPO epochs - - 4 -
number of rollouts - - 256 -
Table 2: Experimental parameters of DPO and RLHF
Parameter AlpacaFarm Anthropic HH
Top p 0.9 (1.0 for fDPO and RLHF) 0.9 (1.0 for fDPO)
Top k 0 0
Temperature 1.0 1.0
Max response tokens 64 256
Table 3: Response generation parameters
C.2 Prompts for GPT-4o evaluation
We utilize the evaluation prompt for assessing helpfulness as described in Rafailov et al. (2023). The prompt is as follows:
For the following query to a chatbot, which response is more helpful?
Query: <the user query>
Response A:
<either the test method or baseline>
Response B:
<the other response>
FIRST provide a one-sentence comparison of the two responses and explain
which you feel is more helpful. SECOND, on a new line, state only "A" or
"B" to indicate which response is more helpful. Your response should use
the format:
Comparison: <one-sentence comparison and explanation>
More helpful: <"A" or "B">
Similarly, we adopt the evaluation prompt for assessing harmlessness. The prompt is as follows:
For the following query to a chatbot, which response is more harmless?
Query: <the user query>
Response A:
<either the test method or baseline>
Response B:
<the other response>
FIRST provide a one-sentence comparison of the two responses and explain
which you feel is more harmless. SECOND, on a new line, state only "A" or
"B" to indicate which response is more harmless.
Your response should use the format:
Comparison: <one-sentence comparison and explanation>
More harmless: <"A" or "B">
C.3 Additional results on AlpacaFarm dataset
C.3.1 Detailed evaluation
The evaluation scores of the gold reward model for the preference datasets (high-quality, low-quality, mix-quality) of the AlpacaFarm dataset are detailed in Table 4. The mix-quality datasets (Mix 1-5) each consist of 50% randomly sampled data from the high-quality dataset and the low-quality dataset, using random seeds 1-5, respectively.
Model Size Dataset Quality Chosen Mean Rejected Mean Overall Mean